# Noé Arnold, Juan Beccuti, Thunj Chantramonklasri, Matthias Hafner, Nicolas Oderbolz Center for Cryptoeconomics

### v1

@@ Abstract
This paper studies how different categories of Ethereum stakers respond to changes in the consensus issuance schedule and how such changes may impact the composition of the staking market. Our findings show that solo stakers are more sensitive to variations in staking rewards compared to ETH holders using centralized exchanges or liquid staking providers. This heightened sensitivity is not explained by the assumptions on cost structure— which would suggest the opposite—but is instead driven by market dynamics. Factors such as superior MEV access and DeFi yields available to other staking methods, as well as the potential presence of inattentive stakers, contribute to this effect. Consequently, a reduction in issuance is likely to crowd out solo stakers while increasing the market share of centralized entities, further concentrating validator power.
I Introduction
As the volume of ETH staked within Ethereum’s Proof-of-Stake (PoS) protocol has grown significantly over the past few years, discussions around the long-term sustainability of the current issuance schedule have become prominent. Central to this debate is the trade-off between the efficient use of validator infrastructure, issuance-driven inflation, and economic security.
A reduction in the consensus issuance schedule has been argued to be important in avoiding an end-game scenario in which most of the circulating supply of ETH is staked. Such an outcome would mean that the supply of staked ETH far exceeds the necessary threshold for economic security, resulting in unnecessary inflationary pressure and inefficient allocation of validator resources. In addition, there are concerns about the potential systemic risks posed by the proliferation of liquid staking tokens and their potential to replace ETH as the de facto currency in the ecosystem. These considerations suggest that it may be prudent to revise the issuance schedule to effectively reduce the incentives for staking.
Conversely, reducing staking rewards may disproportionately impact the profitability of solo stakers, who already represent a small segment of the total validator population. Such a shift could inadvertently accelerate the centralization of Ethereum’s validator ecosystem by driving solo stakers out of the market. In what follows, we aim to test the validity of this argument by examining how different staking methods, and in particular solo stakers, may be affected differently by changes to Ethereum’s consensus issuance policy. Specifically, we focus on answering two key research questions:
  • How do different types of stakers respond to shifts in the issuance schedule?
  • How might these changes impact the decentralization of the Ethereum staking ecosystem?
In the following, we aim to answer these research questions by developing a game-theoretic model of the Ethereum staking market. We run simulations of this model and calibrate it using currently available empirical evidence. Based on the calibrated model, we conduct a comparative analysis of equilibrium outcomes under different consensus issuance schedules.
@@ I Introduction
The empirical literature mostly suggests that solo stakers are more sensitive to changes in consensus issuance yields compared to ETH holders staking through centralized exchanges or liquid staking providers. Consequently, a reduced issuance schedule could be expected to decrease the market share of solo staking while increasing that of centralized exchanges, leading to higher centralization.
@@
Based on our calibrated model, we argue that the cost structures of solo stakers would on their own likely suggest lower supply elasticity with respect to changes in the consensus issuance yield. However, this alone does not fully explain the staking decisions of solo stakers. Instead, their responsiveness is also shaped by market dynamics, where declining issuance intensifies competition for consensus yields. Other staking methods, benefiting from superior MEV access and DeFi yields, are less affected by issuance changes. Strategic solo stakers internalize these competitive pressures, making them more sensitive to shifts in consensus rewards. The results highlights the importance of policy measures like MEV burn in mitigating competitive disadvantages of solo stakers that arise from diminished or highly variable access to MEV revenues.
@@
In addition, we show that reduced issuance disproportionately reduces the profitability of solo staking relative to other staking categories, in part due to the fact that solo staking has fixed costs from the perspective of the ETH holder, while other staking categories do not. While beyond the scope of our model, this reduces the long-term incentives for solo staking and may lead to exits or shifts to solutions such as dSSPs, which can provide stakers with additional DeFi returns and stable MEV revenues.
The structure of this paper is as follows: Section II reviews the existing literature, with a focus on the empirical estimation of the yield elasiticity of staking supply. Section III introduces the game-theoretic model of the Ethereum staking market and outlines its main assumptions and additional model extensions. Section IV reports on the results of the simulation and comparative analysis. Section V discusses the results and Section VI concludes.
II Literature Review
In the following, we focus on research that either estimates the staking behavior of stakers in terms of staking returns or provides empirical evidence on how ETH holders make decisions on which staking method they prefer to use. Several noteworthy contributions have been made to the literature recently.
We turn first to the empirical literature. A recent paper using event studies and empirical elasticity estimation finds that individual stakers in Ethereum exhibit greater sensitivity to relative declines in returns compared to other staking categories (Eloranta and Helminen,, 2025). The authors argue that this suggests that individual solo stakers may be disproportionately affected by changes in network incentives.
@@ II Literature Review
Another recent contribution to the literature is based on survey data and semi-structured interviews (Zhu et al.,, 2025). Here, the authors find that solo stakers exhibit heightened sensitivity to changes in the issuance curve relative to other stakeholders, a finding consistent with previous findings. The authors also find some evidence of market segmentation. Large crypto holders tend to view staking rewards as a secondary benefit, with their primary focus being on ETH price appreciation. These individuals consequently tend to prefer liquid staking solutions. On the other hand, retail investors tend to prioritize convenience and brand trust over short-term price fluctuations or minor yield fluctuations. As a result, they often stake through centralized platforms such as Coinbase and are relatively insensitive to fluctuations in staking returns. Finally, institutional stakers tend to emphasize security and regulatory compliance in their staking decisions, with their primary motivation being to protect their ETH holdings from dilution.
@@
Overall, the available empirical literature provides several key insights. First, solo stakers tend to be more sensitive to fluctuations in staking returns than other staking categories. This finding challenges the conventional wisdom that solo stakers are primarily motivated by altruism and a commitment to Ethereum’s decentralization, rather than financial incentives. Furthermore, the literature suggests that the Ethereum staking market is to some extent segmented based on user preferences. In particular, institutions and retail investors tend to focus on the staking services of centralized exchanges such as Coinbase, while DeFi-native stakers tend to prefer liquid staking solutions such as Lido.
In terms of formally describing the Ethereum staking market, some preliminary research has been presented in which the staking allocation decision of ETH holders across different staking methods is modeled as a function of both monetary returns and non-monetary preferences, such as convenience, trust, and commitment to decentralization (Ma,, 2024). The model predicts that changes in the issuance schedule do not affect an individual’s optimal choice of staking method. Contrary to the empirical evidence, the result is that the total market share of a given staking method remains unchanged when the issuance schedule is adjusted. This result depends on the assumption of linear cost functions and, importantly, does not model strategic interactions among stakers competing for issuance returns. In the following, we extend this research by relaxing the assumption of linear cost functions and by modeling strategic interactions between market participants.
III Theoretical Model
This section presents a game-theoretic model incorporating heterogeneous ETH holders and multiple staking alternatives. It outlines key simplifying assumptions and explores potential model extensions. The model framework facilitates analysis of how differing levels of staker sophistication and SSP fee structures influence staking participation and the overall reward structure.
III-A Baseline Model
@@
Building on the empirical evidence reviewed in Section II, the model differentiates ETH holders based on their preferences and level of technical sophistication, which in turn determines the staking options available to them. Specifically, three distinct types of ETH holders are considered:
@@
Each category consists of a fixed number of players, denoted as $N_{r}$, $N_{t}$, and $N_{ss}$ for Retailers, Techies, and Experts, respectively. For analytical tractability, we assume homogeneity within each group.
In addition, the model includes several staking methods that differ both in the costs they impose on the ETH holder and in the revenue streams they provide. These methods include solo staking, staking through a centralized service provider (cSSP) such as a centralized exchange, and staking through a decentralized service provider (dSSP) that provides additional DeFi revenues through an LST (e.g., Lido).
Consistent with the evidence presented in Section II, we assume a segregated staking market. This means that ETH holders of the Retailer type can either stake with a cSSP or opt out of staking altogether, arguably reflecting their preference for the convenience of centralized exchanges or their limited technical expertise that prevents them from engaging in solo staking or using a dSSP. Techies choose to stake through a dSSP or not to stake at all. This is consistent with Zhu et al., (2025), as they are assumed to value liquidity and prefer decentralized intermediaries that provide additional returns through liquid staking tokens (LSTs). Finally, Experts engage exclusively in solo staking or abstain from staking, reflecting their strong preference for decentralization, which precludes their reliance on any intermediaries. In the following, we will use the terms solo staker and expert interchangeably.
Depending on the staking option they choose, ETH holders may receive revenues from three different sources: 1) consensus layer revenue, 2) MEV revenue, and 3) additional DeFi yield.
Consensus layer revenues are determined by the protocol’s issuance schedule. The annual consensus yield provided to an ETH holder when staking is denoted as $y_{i}(D)$ and under the current issuance schedule is given by
@@
where $D$ represents the total amount of staked ETH staked. In the following, we will essentially compare model results under the current issuance schedule $y_{i}(D)$ with alternative specifications $y_{i}^{\prime}(D)$.
The second revenue stream is MEV revenue. Validators selected as proposers in block production can generate additional revenue from structuring transactions. Let $N=D/32$ denote the total number of validators, the probability that an ETH holder controlling $d_{i}/32$ validators will be selected as a proposer is $d_{i}/D$. Conversely, if a staker joins a staking pool (i.e., stakes with an intermediary), proposers can share MEV revenue with validators. The probability of any validator in the pool being selected as a proposer is $d_{pool}/D$, and a validator with stake $d_{i}$ receives a share of $d_{i}/d_{pool}$. Thus, the expected MEV revenue for a stakeholder in a pool remains the same as for a solo staker. In both cases, the expected annual MEV revenue $y_{v}$ for an ETH holder $i$ is theoretically given by
@@
The third source of income is additional DeFi income. ETH holders using a dSSP can earn additional income by reinvesting the LST they receive. The annual return from this income source is assumed to be exogenous and is simply defined as $y_{d}$.
Staking entails costs for ETH holders, which vary depending on the staking solution. In this model, each staking method is associated with a distinct cost function that may include fixed and nonlinear variable costs

### v2

# Matthias Hafner, Nicolas Oderbolz

@@ Abstract
This paper examines how various categories of Ethereum stakers respond to changes in the consensus issuance schedule, and the potential impact of such changes on the composition of the staking market. To this end, we have develop and calibrate a game-theoretic model of the Ethereum staking market, incorporating strategic interactions between various staking agents. Our findings suggest that solo stakers may be more sensitive to variations in staking rewards than ETH holders using centralized exchanges or liquid staking providers. This increased sensitivity is driven not only by the cost structure of solo staking, but also by the competitive dynamics between different staking solutions in the market. When faced with a downward-sloping issuance schedule, staking agents compete for limited staking yields, and their choice of staking supply affects the revenues of other stakers. Therefore, the presence of other staking methods with access to MEV revenues and other DeFi yield sources when staking, as well as inattentive stakers, puts competitive pressure on solo stakers. Consequently, our model predicts that a reduction in issuance is likely to crowd out solo stakers. We present preliminary empirical evidence to support this result, using an instrumental variable estimation approach to estimate the yield elasticity of staking supply for different staking categories.
1 Introduction
1.1 Context
The consensus issuance schedule, a central element of Ethereum’s proof-of-stake (PoS) consensus protocol, determines how new tokens are created and distributed to validators as a reward for their participation in securing the network. To ensure robust network security, the issuance schedule must be structured to incentivize a sufficient number of validators to participate in staking. At the same time, if the issuance schedule sets rewards too high and facilitates levels of staking that exceed those required for economic security, the resulting inflation will unnecessarily dilute ETH holders. In addition, the resources devoted to running validator nodes could, on the margin, be allocated more efficiently elsewhere. To balance this trade-off, the issuance schedule follows a downward-sloping curve: when the total amount of staked ETH is low, rewards are high, but as more ETH is staked, rewards decrease.
After the introduction of the Proof-of-Stake (PoS) consensus mechanism, the volume of staked ETH steadily increased and, after slightly declining in the past months, currently stands at approximately 35.8 million ETH or around 30% of the total circulating supply. [^1] Data obtained from https://www.stakingrewards.com/asset/ethereum-2-0/analytics, 16.07.2025 As a result of this growth, there has been speculation about potential long-term scenarios in which a significant portion of the circulating supply is locked within the staking protocol. In addition to an unbalanced trade-off between economic security and inflation, as well as an inefficient use of validator infrastructure, this could mean that in the long run, ETH is effectively replaced by more centralized and non-trustless liquid (re-)staking tokens as the predominant form of currency in the ecosystem. Motivated by these concerns, several proposals aimed at modifying the issuance schedule to effectively reduce staking incentives set by the protocol have emerged.(Drake,, 2024; Elowsson, 2024a, ; Elowsson, 2024b, ; Elowsson, 2024c, )
However, a change in the consensus issuance schedule could affect different types of stakers in different ways, which could have significant consequences for the composition of the Ethereum validator set, and therefore the degree of decentralization of the protocol. There exist different staking solutions in the Ethereum staking market, each with distinct implications for the protocol’s decentralization (Kotelskiy et al.,, 2024). While individual solo stakers contribute significantly to decentralization, centralized exchanges control substantial portions of staked ETH within proprietary infrastructures, introducing centralization risks. Liquid staking solutions, like Lido, tend to occupy an intermediate position: they typically feature decentralized governance structures and distribute their stake across multiple uncorrelated staking service providers. Liquid staking tokens benefit from strong network effects, which can enable a single issuer to gain significant market share and, consequently, substantial influence within the Ethereum ecosystem (Neuder,, 2023).
Some argue that reducing the staking rewards issued by the protocol would better balance the trade-off between economic security and inflation. It would also limit the continued adoption of liquid staking tokens (LSTs) and thus reduce the influence of liquid staking protocols. However, opponents maintain that Ethereum’s issuance policy should primarily serve to maximize the network’s economic security. They caution that reducing staking rewards could push solo stakers and smaller operators, who face higher setup costs and lack the economies of scale enjoyed by larger operators, out of the market. Consequently, staking could become more centralized if consensus issuance is reduced, which would ultimately weaken the protocol’s economic security and long-term stability (Kotelskiy et al.,, 2024).
1.2 Research Questions and Methodology
The present research aims to contribute to the above discussion by developing a formal understanding of how different players in the Ethereum staking market form staking decisions and how these may be affected by changes in the issuance schedule. Specifically, two main research questions are addressed sequentially:
  1. 1.
    How can we characterize the staking supply from different categories of stakers?
  2. 2.
    How does a change in the issuance schedule affect the staking supply of these stakers?
To address these research questions, we review the relevant existing literature and provide a short overview of the staking market. We then develop a formal model describing the Ethereum staking market equilibrium in terms of the costs and revenues faced by individual agents within the ecosystem. By conducting numerical simulations of this model, we aim to understand the factors that drive staking behavior and consequently equilibrium staking outcomes. Additionally, we calibrate the model using specific parameter values that try to approximate the current market environment, evaluating how specific changes in the issuance schedule may affect staking market equilibrium.
1.3 Summary of Main Findings
@@
Using a combination of empirical evidence and a calibrated theoretical model, we demonstrate that solo stakers may be more susceptible to fluctuations in consensus issuance yields than ETH holders who stake through centralized exchanges or liquid staking providers. Our findings suggest that a reduced issuance schedule could lead to a reduced market share of solo staking and an increase in the market share of centralized exchanges, leading to greater centralization.
@@
In our model, we begin by assuming that the cost structures of solo stakers alone would suggest lower supply elasticity in response to changes in the consensus issuance yield, compared to other staking methods. Even so, we find that other factors influence the staking decisions of solo stakers within the market, on net resulting in the staking supply of solo stakers being relatively sensitive to changes in the issuance schedule. Importantly, their staking supply decision is also shaped by market dynamics. Given a downward-sloping issuance schedule, stakers compete for issuance yields. A reduced issuance intensifies this competition further. Since other staking methods benefit from superior MEV access and DeFi yields, they are less affected by the change in issuance. This creates a crowding-out effect for solo stakers. Strategic solo stakers internalize these competitive pressures, making them more sensitive to shifts in consensus rewards.
@@
Our results emphasize the importance of policy measures such as MEV burn in mitigating the negative impact of issuance reductions on solo staking. Moreover, our model predicts that reduced issuance disproportionately reduces the profitability of solo staking compared to other staking categories. While beyond the scope of our model, this may lower long-term incentives for solo staking and may drive exits or shifts in supply towards liquid staking solutions, which can offer stakers additional DeFi yields.
1.4 Structure
This paper is structured as follows: Section 2 reviews the relevant existing literature. Section 3 provides a brief data-driven overview of the Ethereum staking market. Section 4 introduces a game-theoretic model of the Ethereum staking and outlines its main assumptions. Additionally, this section reports on the numerical simulation of the model, as well as the comparitive analysis under different issuance schedules. In Section 5, we estimate the sensitivity of the staking supply of different staking categories to changes in staking returns using an instrumental variable approach. Section 6 discusses the results presented in the previous sections and Section 7 concludes.
2 Literature Review
First, Eloranta and Helminen, (2025) conduct an empirical analysis of the Ethereum staking market in view of a potential change in the consensus issuance schedule. They identify a significant relationship between staking operation size and variations in staking returns, with historical data indicating that larger staking pools consistently achieve higher returns compared to single-validator pools. Moreover, they find that solo stakers demonstrate greater sensitivity to relative declines in yield compared to other staking categories. They suggest that this heightened responsiveness may indicate that individual validators would be disproportionately impacted by changes in network incentives.
@@ 2 Literature Review
Zhu et al., (2025) conduct interviews with various stakeholder groups in the Ethereum staking market. Based on these interviews, they argue that solo stakers exhibit heightened sensitivity to changes in the issuance curve relative to other stakeholders, a finding consistent with Eloranta and Helminen, (2025). Moreover, the authors find that large crypto holders generally regard staking rewards as a secondary benefit, with their primary focus being on ETH price appreciation. These individuals also tend to prefer liquid staking solutions for the utility of their LSTs provide in DeFi. Additionally, the study argues that retail investors tend to prioritize convenience and brand trust over short-term price fluctuations or minor yield variations. As a result, they often stake through platforms like Coinbase and are relatively insensitive to fluctuation in staking yields. Lastly, institutional stakeholders tend to emphasize security and regulatory compliance in their staking decisions, with their primary motivation being the preservation of their ETH holdings against dilution.
@@
Overall, the available empirical literature recovers several key insights. First, solo stakers are consistently found to demonstrate greater sensitivity to fluctuations in staking yields compared to other staking categories. This finding challenges the conventional assumption that solo stakers are primarily motivated by altruism and a commitment to Ethereum’s decentralization, rather than financial incentives. Furthermore, the literature suggests that the Ethereum staking market is to some degree segmented along different staking solutions. In particular, institutions and retail investors tend to concentrate around the staking services of centralized exchanges, such as Coinbase, while DeFi-native stakers tend to prefer liquid staking solutions like Lido.
Regarding the modeling approach adopted in this study, perhaps the most relevant existing work is that of Ma, (2024), who models the staking allocation process of ETH holders across different staking methods as a function of both monetary returns and non-monetary preferences, such as convenience, trust, and commitment to decentralization. Under the assumption of linear cost functions, the model predicts that changes in the issuance schedule do not affect an individual’s optimal choice of staking method. The result is that the overall market share of a given staking method remains unchanged when the issuance schedule is adjusted. While this modeling approach represents an important first step in modeling the distribution of staking supply in relation to issuance levels and the costs associated with different forms of staking, the results run counter to the empirical findings above. In particular, the observed higher sensitivity of solo stakers to changes in staking yields suggests that their market share would likely decline in response to reductions in issuance. Consequently, relaxing the assumption of linear cost functions may be important to achieving better alignment with the empirical evidence. Among other novel features, the theoretical model developed in the following takes a step in this direction by allowing for non-linear staking cost structures.
3 Historical Staking Supply
To provide an overview of the Ethereum staking market, we illustrate data on the overall market shares of different staking methods [^2] Source: Dune query by user hildobby, see https://dune.com/queries/1941407/3202651. and on the exits and entries of validators within different staking methods [^3] Source: Rated Network through MVI Grants Program between 1 December 2020 and 27 May 2024. In particular, Figures 1 and 2 illustrate the evolution of staked ETH across different staking categories over time, and Figure 3 shows the entry and exit dynamics of validators across these solutions. While a portion of staking supply cannot be reliably attributed to a specific staking solution, thus introducing some degree of uncertainty to any conclusions drawn from the data, several distinct patterns emerge.
First, both centralized exchanges and liquid staking solutions have grown significantly over time, collectively capturing a substantial market share. More recently, however, the rapid rise of liquid restaking has coincided with a decline in staking via centralized exchanges and liquid staking solutions, both in absolute terms and as a share of total staked ETH. This trend suggests an element of competition between different staking methods. While large liquid staking providers benefit from strong network effects due to the issuance of LSTs, these effects seem not to have been sufficient to fully insulate them from market share erosion following the introduction of new staking innovations.
Additionally, solo staking has experienced a steady decline, both in absolute terms and as a proportion of total staking. As shown in Figure 3, solo staking saw significant validator entry in the early stages of Ethereum’s PoS transition. However, this trend reversed over time, with validator exits outweighing new entries after the Shapella upgrade, which enabled stakers to withdraw their staked ETH. The Shapella upgrade also represents a noticeable shift in user behavior for staking via centralized exchanges and via liquid staking solutions. Both categories saw an increase in validator entries, while only centralized exchanges saw a sharp increase in exits. On this basis, it appears that users of centralized exchanges also valued the improved liquidity introduced by the Shapella upgrade. Of course, if this is the case, it begs the question of why users of centralized exchanges ever chose these solutions over liquid staking, which by design allowed users to retain the liquidity of their stake even before the Shapella upgrade. It may therefore be reasonable to infer that there are other user preferences that segment the staking market to a degree that has allowed centralized exchanges to retain a substantial market share even during periods when liquid staking solutions could offer far greater liquidity.
Taken together, this supports the findings of Zhu et al., (2025), that the Ethereum staking market exhibits a degree of segmentation based on the type of staking solution and user preferences. While competition certainly appears to exist between liquid staking and liquid restaking—implying a shared customer base—the broader dynamics of staking supply seem to suggest that solo staking, liquid staking, and staking via centralized exchanges cater to distinct user preferences and constraints.
Figure 1: Staked ETH by staking method
Source: Dune query by user hildobby, see https://dune.com/queries/1941407/3202651.
Figure 2: Share of staked ETH by staking method
Source: Dune query by user hildobby, see https://dune.com/queries/1941407/3202651
Figure 3: Validator entry and exit over time
((a)) Total
((b)) Solo Staking
((c)) CEX
((d)) Liquid Staking
Source: Rated Network
4 Theoretical Model
The following section presents a game-theoretic framework through which we can model the strategic staking decisions of various market participants in relation to the Ethereum issuance schedule, alternative revenue sources and staking costs. In particular, the model framework enables us to analyze how differing levels of staker sophistication and SSP fee structures influence staking participation and the overall reward structure. For an overview of the parameters defined by the model, please refer to Table 2.
4.1 Agent Types
@@
Building on the empirical findings in Section 3, the model differentiates ETH holders based on their preferences and level of technical sophistication, which in turn determines the staking options available to them. Specifically, three distinct types of ETH holders are considered:
@@
Each category comprises a fixed number of stakers, denoted as $N_{r}$, $N_{t}$, and $N_{ss}$ for Retailers, Techies, and Experts (a category composed entirely of solo stakers as we explain below), respectively. For analytical tractability, we assume homogeneity within each group.
4.2 Segmented Staking Market
In addition, the model incorporates several staking methods that differ in the costs they impose on ETH holders and the revenue streams they provide. These methods include solo staking; staking through a centralized service provider (cSSP), such as a centralized exchange or a professional direct staking delegation company like Kiln; and staking through a decentralized service provider (dSSP) that provides additional DeFi yields through an LST, such as Lido. We do not model restaking explicitly.
Motivated by the market dynamics described in Section 3 and the evidence presented by Zhu et al., (2025), we model the preferences of the different agent types to depict a segmented staking market. [^4] Segmented staking markets are also considered in other studies, such as Kotelskiy et al., (2024). That is, each ETH holder type selects a single staking method, and does not diversify across options. Specifically:
  • Assuming limited technical ability and a preference for ease of use, retailers either stake with a cSSP or abstain, which prevents them from engaging in solo staking or using a dSSP.
  • Techies prefer dSSPs over cSSPs because they receive higher returns from LSTs. That is, we model revenue and cost functions such that Techies’ profits are larger than those of Retailers. As such, techies are assumed to value liquidity and favor decentralized intermediaries that provide additional yields through liquid staking tokens (LSTs).
  • Experts are assumed to exclusively choose to solo stake—despite potentially lower profits—due to an intrinsic (but unmodeled) utility derived from contributing to decentralization. As a consequence, Experts solo stake or leave the staking market.
In sum, each ETH holder allocates their entire stake to the option that yields the highest utility for their type. This framework emphasizes clear segmentation rather than hybrid strategies. [^5] An alternative approach could involve portfolio management. However, this kind of approach makes more sense when the different asset risk-adjusted returns have low or negative correlation. This might not be the case when considering only different kind of staking methods as in our model.
4.3 Staking Revenues
Depending on their choice of staking option, ETH holders may receive revenues from three different sources: 1) consensus layer revenue, 2) MEV revenue, and 3) additional DeFi yield.
The first is consensus layer revenue, determined by the protocol’s issuance schedule. The annual consensus yield provided to an ETH holder when staking is denoted as $y_{i}(D)$. Under the current issuance schedule, the annual consensus yield is given by
@@ 4.3 Staking Revenues
where $D$ represents the total amount of ETH staked in the protocol. In the following, we will essentially compare model outcomes under the current issuance schedule $y_{i}(D)$ to alternative specifications $y_{i}^{\prime}(D)$.
The second revenue source is MEV revenue. Validators selected as proposers in block production can extract additional revenue from structuring transactions, referred to in the following as MEV revenue. Letting $N=D/32$ denote the total number of validators, the probability that an ETH holder controlling $d_{i}/32$ validators is chosen as a proposer is $d_{i}/D$. Conversely, if a staker joins a staking pool (i.e. stakes with an intermediary), proposers may share MEV revenues with attesters. The probability that the proposer belongs to the pool is $d_{pool}/D$, and a validator in the pool with stake $d_{i}$ receives a share of $d_{i}/d_{pool}$. Thus, the expected MEV revenue for a staker in a pool remains the same as for a solo staker. In both cases the expected annual MEV revenue $y_{v}$ for an ETH holder is theoretically given by
@@
The third source of revenue is additional DeFi yield. ETH holders utilizing a dSSP may earn additional yields by reinvesting the LST they receive. The annual yield from this source of revenue is simply defined as $y_{d}$.
4.4 Staking Costs
From the perspective of the ETH holder, staking is also associated with costs. In the present model, the cost function faced by an ETH holder is allowed to differ for each staking solution. We define the cost function to include fixed costs and non-linear variable costs: